Find the volume generated by revolving the area bounded by y=e^x, x=1, x=0, about y-axis. Sketch the curve indicating the generating area.
1
Expert's answer
2020-12-21T17:00:06-0500
"y=e^x, x=0, x=1"

"V=\displaystyle\int_{0}^12\pi x(e^x)dx"

"\displaystyle\int xe^xdx"


"\int udv=uv-\int vdu"

"u=x, du=dx"

"dv=e^xdx, v=\displaystyle\int e^xdx=e^x"


"\displaystyle\int xe^xdx=xe^x-\displaystyle\int e^xdx=xe^x-e^x+C"

"V=\displaystyle\int_{0}^12\pi x(e^x)dx=2\pi\big[xe^x-e^x\big]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=2\pi(e-e-(0-1))=2\pi (units^3)"

"V=2\pi\ cubic\ units"